In a close private economy the level of planned investment is 300 and the savings function is S=-100 +0.25Y. The equilibrium level of income and consumption respectively is
1
Expert's answer
2020-06-16T12:20:29-0400

"S=-100+0.25Y"

"I=300 (investment)"

Output in equilibrium "Y=C+I+G+X-M"

Consumption function

"C=a+b(Y-T)"

Saving function

"S=-a+(1-b)(Y-T)"

"Y=100+0.75Y+300"

"Y=400+0.75Y"

"C=100+0.75Y"